OpenStudy (anonymous):

Differentiate. y = 6x/(9 − tan x) Any help would be greatly appreciated!

OpenStudy (aum):

\[ \left( \frac{f}{g} \right )' = \frac{gf'-fg'}{g^2} \]

OpenStudy (anonymous):

I got an answer but it says its wrong:/ I got (6(9-tans-6xsec^2x))/(9-tans)^2. Can you tell what I did wrong?

OpenStudy (anonymous):

@aum

OpenStudy (aum):

\[ \left( \frac{6x}{9-\tan(x)} \right )' = \frac{(9-\tan(x))(6x)'-(6x)(9-\tan(x))'}{(9-\tan(x))^2} = \\ \frac{6(9-\tan(x)) - 6x(-\sec^2(x))}{(9-\tan(x))^2} = \frac{6(9-\tan(x) + x\sec^2(x))}{(9-\tan(x))^2} \]

OpenStudy (anonymous):

THANK YOU SO MUCH @aum

OpenStudy (aum):

You are welcome.